Circular No. 6853

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980329 F. Frontera, Istituto Tecnologie e Studio Radiazioni Extraterrestri,

Bologna; E. Costa, L. Piro, and P. Soffitta, Istituto Astrofisica Spaziale, CNR, Rome; J. in ’t Zand, Space Research Organization of the Netherlands (SRON), Utrecht; L. Di Ciolo and A. Tesseri, BeppoSAX Science Operation Center (SOC), Rome; and M. Smith, SOC and SRON report: “The BeppoSAX Gamma Ray Burst Monitor was triggered by a γ-ray burst on Mar. 29.1559 UT, corresponding to BATSE trigger 6665. A preliminary analysis shows a duration of 15 s and a peak intensity of 6000 counts/s (40–700 keV). This is the brightest burst detected so far by the BeppoSAX GRBM simultaneously with the Wide Field Cameras (WFC). The peak intensity measured in the WFC is ∼ 6 Crab (2–26 keV). The source position is α = 7^{h}02^{m}41^{s}, δ = +38^{o}50′.7 (equinox 2000.0; error radius 3′). A BeppoSAX follow-up observation is in progress.”

AX J0051–73.3 AND AX J0058–72.0 J. Yokogawa and K. Koyama, Kyoto University, report: “ASCA de-

tected two new x-ray pulsars in the Small Magellanic Cloud, designated AX J0051–73.3 and AX J0058–72.0. AX J0051–73.3 was detected on 1997 Nov. 13 at α = 0^{h}50^{m}47^{s}, δ = −73^{o}15′51′′ (equinox 2000.0; ± 1′ at 90-percent confidence), with a pulse period of 323.2 ± 0.4 s and a quasi-sinusoidal pulse shape. Its spectrum was well fitted with a power law having a photon index of 0.6 ± 0.2; the x-ray flux in the band 0.7–10.0 keV was 3.6 × 10^{−12} erg s^{−1} cm^{−2}. AX J0058–72.0 was detected on 1997 Nov. 14 at α = 0^{h}57^{m}57^{s}, δ = −72^{o}02′46′′ (equinox 2000.0; ± 1′ at 90-percent confidence). Its pulse period was 280.4 ± 0.3 s, with a single-peaked pulse shape; its spectrum was well fitted with a power law having a photon index of 1.0 ± 0.3, and its x-ray flux in the band 0.7–10.0 keV was 2.4 × 10^{−12} erg s^{−1} cm^{−2}.”

SUPERNOVAE 1997dq AND 1998W S. Moretti and S. Tomaselli, Forlí, Italy, report the following unfiltered

CCD magnitude estimates using GSC reference stars: SN 1997dq in NGC 3810 (cf. IAUC 6770), Mar. 26.92 UT, 17.6; SN 1998W in NGC 3075 (cf. IAUC 6843), Mar. 26.91, 17.1.

1998 March 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green